OpenStudy (anonymous):

how do you do 9x^2 -81=0

hero (hero):

(3x-9)(3x+9) = 0

OpenStudy (anonymous):

how do you dot that? i never learned

hero (hero):

It's called difference of squares

hero (hero):

If you see nothing but squares and then they are being subtracted, you just use this rule: ax^2 - b^2 = (ax+b)(ax-b)